book JET Extension Tags Bradford Matthew Arant Sr. document Introduction JET will pass through the untagged areas to the output without any modifications. Data contained

within the tags may modify their containers before placing any output. The space taken by the tag itself is not passed to the output and will not appear in the output.

Skip Blank Lines options on containers will skip passing any blank lines or line containing only whitespace to the output.

Tags and Attributes XML style tagging is used to insert functions and other behaviors into the script. As the document

contents are output a tag may be inserted to perform the desired function and affect the output. Tag attributes are used to pass parameters into the operations and functionality for each tag. Outputting a database from mysql may be performed using the mysqltagincombinationwiththeifrow, whilerowandsqltags. User Defined Tags and Tag Liraries You can define your own tag definitions and use them within your JET scripts. Keywords are not defined as part of the tag definition and any attribute added to the implementation

of the tag is passed into the tag process as a local variable for that container.

Tag definitions can also be defined as scoped or unscoped, as follows: scoped - local variables using the #[n] format can be scoped to just your tag contents. Access to local

variables where your tag is being used must use the ’parent’ scope option to retrieve variables from the parent.

unscoped - local variables are shared with the parent tag environment and are not unique to your tag definition.

A ¡container /¿ tag can be inserted into the container and will take the contents of the container at runtime and insert it into the output when processing the user defined tag definition. The container tag of the the ¡container /¿ is processed before the container containing the ¡container /¿ so that local variabls can be set and processed that will affect the tags overall output.

Tags are many times nested with tags running within the container of a higher tag. Variables and Variable Types Variables are used to contain dynamic content values and can be sourced from several locations and

limited in scope depending on use. Some variable types can be set to reflect script state or are read only from other outside sources of data.

Jet tags can be either scoped or unscoped. The following is a list of variable types and a brief description of their source: itemize

G lobal - values can be initiaized using the set tag and specifying the scope of global. L ocal - K eyword - within a tag’s container we may access the keyword values specified on the containing tag.